OpenStudy (anonymous):

K, can someone check my work I'm trying to find the surface area of a polar coordinate system

OpenStudy (anonymous):

Problem r=sin2theta

OpenStudy (anonymous):

\[1/2\int\limits_{0}^{\pi/2} \sin^22\theta \]

OpenStudy (anonymous):

\[1/2\int\limits_{0}^{\pi/2} (1-\cos4\theta)d \theta\]

OpenStudy (anonymous):

\[1/2(\theta - 1/2 \sin4\theta)\]

OpenStudy (anonymous):

\[1/2(\pi/2-1/2\sin4(\pi/2) - 0 \]

OpenStudy (anonymous):

= pi/4 ? does everything look good

OpenStudy (anonymous):

hmmm... i keep getting pi/8 for the area of that 1 leaf in the first quadrant....

OpenStudy (anonymous):

i don't beleive you u-sub'd correctly. you have: \[\huge \int\limits \cos(4\theta) d \theta\] but that should be \[\huge \int\limits {1 \over 4}\sin(4 \theta)\] right?

OpenStudy (anonymous):

thats right my mistake thanks!

OpenStudy (anonymous):

my bad, i typo'd that. Should just be 1/4*sin(4x), no integral sign. My bad.